OpenStudy (anonymous):

Use synthetic division to divide (5x^3+18x^2+7x-6) divided by (x+3). Please show steps if you can so I can try to understand whats going on :)

OpenStudy (anonymous):

I got 5x^2+3x-2, would that be the final answer or do I need to do more?

OpenStudy (anonymous):

You need to change it to a polynomial \[5x^ {2} + 3x - 2\] Can be changed to (5x-2) (x+1)

OpenStudy (whpalmer4):

Let's check your answer: \[(x+3)(5x^2+3x-2) = 5x^3+3x^2-2x+15x^2+9x-6 \]\[= 5x^3+18x^2+7x-6\] That's all you need to do to divide it, I don't see any requirement that you need to factor the quotient.

OpenStudy (anonymous):

then 5x-2=0 and x+1=0 so x=2/5,-1

OpenStudy (anonymous):

that should be the final answer

OpenStudy (whpalmer4):

@QWERTYCraft where are you getting this notion that she's supposed to find the roots?

OpenStudy (anonymous):

Most teacher usually ask for them, i'm just trying to show it, just in case they ask for it

OpenStudy (anonymous):

thank you both!